# Topolow: Force-Directed Euclidean Embedding of Dissimilarity Data with Robustness Against Non-Metricity and Sparsity

### v1

@@ Abstract
The problem of embedding a set of objects into a low-dimensional Euclidean space based on a matrix of pairwise dissimilarities is fundamental in data analysis, machine learning, and statistics. However, the assumptions of many standard analytical methods are violated when the input dissimilarities fail to satisfy metric or Euclidean axioms. We present the mathematical and statistical foundations of Topolow, a physics-inspired, gradient-free optimization framework for such embedding problems. Topolow is conceptually related to force-directed graph drawing algorithms but is fundamentally distinguished by its goal of quantitative metric reconstruction. It models objects as particles in a physical system, and its novel optimization scheme proceeds through sequential, stochastic pairwise interactions, which circumvents the need to compute a global gradient and provides robustness against convergence to local optima, especially for sparse data. Topolow maximizes the likelihood under a Laplace error model, robust to outliers and heterogeneous errors, and properly handles censored data. Crucially, Topolow does not require the input dissimilarities to be metric, making it a robust solution for embedding non-metric measurements into a valid Euclidean space, thereby enabling the use of standard analytical tools. We demonstrate the superior performance of Topolow compared to standard Multidimensional Scaling (MDS) methods in reconstructing the geometry of sparse and non-Euclidean data. This paper formalizes the algorithm, first introduced as Topolow in the context of antigenic mapping in Arhami and Rohani, (2025) (open access), with emphasis on its metric embedding and mathematical properties for a broader audience. The general-purpose function Euclidify is available in the R package topolow.
@@ 1 Introduction
The task of representing the similarity or dissimilarity of objects in a low-dimensional Euclidean space is a cornerstone of data analysis. These objects can range from abstract concepts, such as customer perception of different brands (Cooper,, 1983; Borg et al.,, 2018), to complex biological measurements such as protein folding (Holm and Sander,, 1996), or information constructs that do not necessarily satisfy the axioms of a Euclidean space and, hence, are incompatible with many standard statistical or machine learning tools (Duin,, 2005). When dissimilarity data violate metric axioms, they cannot be embedded in a vector space where linear algebraic operations are valid. The fundamental issue is that coordinate-based methods assume data points exist in a vector space equipped with a valid inner product structure. This inner product ⟨·,·⟩ –used in many models, from linear regression to artificial intelligence systems– induces a metric satisfying all axioms (Burago et al.,, 2001).
@@
1.1 The Challenge of Non-Euclidean Dissimilarities
The fundamental challenge of confronting and resolving non-Euclidean structures has deep roots in the history of geometry. The logical foundations of non-Euclidean space were first systematically investigated by the 11th-century Persian mathematician Omar Khayyam in his treatise ”The Difficulties in the Postulates of Euclid’s Elements(Khayyam,, 1936; Struik,, 1958). Many standard data analysis techniques, particularly those from linear algebra, are fundamentally incompatible with non-Euclidean data.
@@ 1.1 The Challenge of Non-Euclidean Dissimilarities
Similarly, clustering algorithms are affected by non-Euclidean dissimilarities. Centroid-based methods, such as $k$-means, become unusable as they require coordinates to compute means. Distance-based methods like hierarchical clustering will run, but can produce counter-intuitive results if the data violate the triangle inequality ($D_{ik}>D_{ij}+D_{jk}$), a key metric property. Such violations disrupt the fundamental notion of ”closeness” upon which these algorithms depend.
Therefore, for a vast significant range of common data, a critical preliminary step is to find a faithful Euclidean representation. An embedding algorithm that can robustly map non-Euclidean dissimilarities into a valid Euclidean space acts as an essential bridge, making the data amenable to the entire suite of standard analytical tools. Topolow is designed to be this bridge.
@@ 2.4.1 Likelihood for Censored Data
Real-world data sometimes contain censored measurements (e.g.,values below a detection limit, $c^{L}$, recorded as ‘$<$$c^{L}$‘, or above a saturation limit, , $c^{R}$, recorded as ‘$>$$c^{R}$‘). Topolow incorporates these directly into the likelihood function.
@@
[…]
@@ 3.1.1 Data Generation
We generated a series of synthetic dissimilarity matrices with known challenging properties. The process begins by creating a set of ground-truth coordinates for $m=50$ objects in a high-dimensional space ($\mathbb{R}^{10}$) with a predefined cluster structure. From these coordinates, a foundational Euclidean distance matrix is computed and systematically distorted through a series of non-linear transformations, including range-dependent power transformations and the addition of asymmetric noise. This procedure yields a complete but non-Euclidean dissimilarity matrix, $M$, which serves as the ground truth for our evaluations. Then, we randomly removed 30% of the pairwise dissimilarities to introduce missing values and used the resulting matrix ($D$) as the input of the algorithms.
@@
    Multi-Scale Analysis: For a fixed sparsity level (30% missing), we generated three variants of $D$ with $m=25,50,$ and 100100100 objects.
@@ 3.1.3 Evaluation Metrics
Performance was assessed by comparing the distances in the resulting Euclidean embedding, $\hat{D}$, to the complete, ground-truth non-Euclidean matrix, $M$.
@@ 3.2 Data Characteristics
A critical first step was to quantify the non-Euclidean nature of our synthetic data. By computing the Gram matrix and its eigenvalues, we can measure the deviation from Euclidean geometry. The Non-Euclidean Deviation Score, calculated as the ratio of the sum of the absolute values of negative eigenvalues to the sum of positive eigenvalues (Gower,, 1985), provides a standardized measure of this distortion. For our primary test case ($m=50$, 30% missing), the deviation score was 0.52, with the negative eigenvalues accounting for over a third of the total variance. This confirms the dataset is non-Euclidean, presenting a significant challenge for any embedding algorithm.
@@ 3.3.1 Accuracy and Distance Reconstruction
In the main test case ($m=50$, 30% sparsity), Topolow demonstrated substantially higher accuracy in reconstructing the ground-truth dissimilarities compared to both Classical and Iterative MDS. As shown in Table 1, Topolow achieved a mean Normalized Stress of 0.190±0.00030.190$\pm$ 0.00030.190 ± 0.0003, an improvement over Classical MDS (0.602) and Iterative MDS (0.552).
Table 1: Performance comparison on a non-Euclidean dataset of 50 objects with 30% missing values. Values for stochastic methods are mean $\pm$ s.d. over 50 runs. Lower stress is better; higher $R^{2}$ is better.
@@
Figure 1: Shepard plots comparing embedded Euclidean distances to original non-Euclidean dissimilarities for the best run of each method ($m=50$, 30% missing). The identity line (dashed) represents perfect reconstruction. Topolow’s embedding (right) shows a strong linear relationship and low distortion. Classical MDS (left) and Iterative MDS (center) show substantially more scatter and deviation from the ideal.
@@ 3.3.2 Robustness to Data Sparsity
At 90% sparsity—an extreme condition where only 10% of dissimilarities are known—Topolow achieved a mean Normalized Stress of 0.5100.5100.510. This result is particularly striking as it is better than the performance of both MDS methods at the much lower 30% sparsity level. This demonstrates the power of Topolow’s approach: the interconnected system of springs enforces a global geometric consistency that allows for robust inference of the overall structure, even when the vast majority of direct measurements are absent. This is fundamentally more effective than attempting to embed a matrix where 90% of the values have been imputed.
@@ 4.1 The Input Dissimilarity Space
[…]
@@ 5 Conclusion
We have presented the mathematics and evaluations of Topolow, an algorithm for Euclidean embedding of potentially sparse, noisy, and non-metric dissimilarity data. By adapting concepts from force-directed graph drawing for the purpose of quantitative metric reconstruction and employing a gradient-free, sequential pairwise optimization, Topolow overcomes the primary limitations of MDS methods. Its foundation in a Laplace error model provides robustness to outliers and censored data, and its integrated, likelihood-based approach to dimensionality selection removes the need for ad-hoc choices. Our empirical evaluation on severely non-Euclidean synthetic data demonstrates that Topolow consistently produces more accurate and stable embeddings than standard MDS methods, particularly as data sparsity and problem size increase. This theoretical robustness translates to practical advantages, as shown in its original application (Arhami and Rohani,, 2025), where Topolow achieved comparable accuracy to MDS for well-structured data of antigenic dissimilarities of influenza viruses but demonstrated a 56% and 41% improvement for more complex dengue and HIV data, respectively. Furthermore, it exhibited orders of magnitude greater run-to-run stability and, unlike MDS, consistently produced a complete embedding of all viruses, a critical feature for sparse datasets. The algorithm’s ability to operate on general, non-metric dissimilarity data is one of its most critical features, providing a necessary bridge between raw, real-world measurements and the vast suite of analytical tools that require valid Euclidean inputs. This makes Topolow a powerful and widely applicable tool for data analysis, visualization, and dimensionality reduction.

### v2

@@ Abstract
The problem of embedding a set of objects into a low-dimensional Euclidean space based on a matrix of pairwise dissimilarities is fundamental in data analysis, machine learning, and statistics. However, the assumptions of many standard analytical methods are violated when the input dissimilarities fail to satisfy metric or Euclidean axioms. We present the mathematical and statistical foundations of Topolow, a physics-inspired, gradient-free optimization framework for such embedding problems. Topolow is conceptually related to force-directed graph drawing algorithms but is fundamentally distinguished by its goal of quantitative metric reconstruction. It models objects as particles in a physical system, and its novel optimization scheme proceeds through sequential, stochastic pairwise interactions, which circumvents the need to compute a global gradient and provides robustness against convergence to local optima, especially for sparse data. Topolow maximizes the likelihood under a Laplace error model, robust to outliers and heterogeneous errors, and properly handles censored data. Crucially, Topolow does not require the input dissimilarities to be metric, making it a robust solution for embedding non-metric measurements into a valid Euclidean space, thereby enabling the use of standard analytical tools. We demonstrate the superior performance of Topolow compared to standard Multidimensional Scaling (MDS) methods in reconstructing the geometry of sparse and non-metric data. This paper formalizes the algorithm, first introduced as Topolow in the context of antigenic mapping in Arhami and Rohani, (2025) (open access), with emphasis on its Euclidean embedding and mathematical properties for a broader audience. The general-purpose function Euclidify is available in the R package topolow.
@@ 1 Introduction
The task of representing the similarity or dissimilarity of objects in a low-dimensional Euclidean space is a cornerstone of data analysis. These objects can range from abstract concepts, such as customer perception of different brands (Cooper,, 1983; Borg et al.,, 2018), to complex biological measurements such as protein folding (Holm and Sander,, 1996), or information constructs that do not necessarily satisfy the metric axioms and, hence, are incompatible with many standard statistical or machine learning tools (Duin,, 2005). When dissimilarity data violate metric axioms, they cannot be embedded in a vector space where linear algebraic operations are valid. The fundamental issue is that coordinate-based methods assume data points exist in a vector space equipped with a valid inner product structure. This inner product ⟨·,·⟩ –used in many models, from linear regression to artificial intelligence systems– induces a metric satisfying all axioms (Burago et al.,, 2001).
@@
1.1 The Challenge of non-metric Dissimilarities
The fundamental challenge of confronting and resolving non-metric structures has deep roots in the history of geometry. The logical foundations of non-Euclidean space were first systematically investigated by the 11th-century Persian mathematician Omar Khayyam in his treatise ”The Difficulties in the Postulates of Euclid’s Elements(Khayyam,, 1936; Struik,, 1958). Many standard data analysis techniques, particularly those from linear algebra, are fundamentally incompatible with non-metric data.
@@ 1.1 The Challenge of non-metric Dissimilarities
Similarly, clustering algorithms are affected by non-metric dissimilarities. Centroid-based methods, such as $k$-means, become unusable as they require coordinates to compute means. Distance-based methods like hierarchical clustering will run, but can produce counter-intuitive results if the data violate the triangle inequality ($D_{ik}>D_{ij}+D_{jk}$), a key metric property. Such violations disrupt the fundamental notion of ”closeness” upon which these algorithms depend.
Therefore, for a vast significant range of common data, a critical preliminary step is to find a faithful Euclidean representation. An embedding algorithm that can robustly map non-metric dissimilarities into a valid Euclidean space acts as an essential bridge, making the data amenable to the entire suite of standard analytical tools. Topolow is designed to be this bridge.
@@ 2.4.1 Likelihood for Censored Data
Real-world data sometimes contain censored measurements (e.g.,values below a detection limit, $c^{L}$, recorded as $c^{L}$‘, or above a saturation limit, , $c^{R}$, recorded as $c^{R}$‘). Topolow incorporates these directly into the likelihood function.
@@
For a Laplace($0,b$) distribution, the CDF is $F(z)=\frac{1}{2}\exp(z/b)$ for $z<0$, and $F(z)=1-\frac{1}{2}\exp(-z/b)$ for $z\geq 0$. The survivor function is S(z)=1F(z)S(z)=1-F(z).
@@ 3.1.1 Data Generation
We generated a series of synthetic dissimilarity matrices with known challenging properties. The process begins by creating a set of ground-truth coordinates for $m=50$ objects in a high-dimensional space ($\mathbb{R}^{10}$) with a predefined cluster structure. From these coordinates, a foundational Euclidean distance matrix is computed and systematically distorted through a series of non-linear transformations, including range-dependent power transformations and the addition of asymmetric noise. This procedure yields a complete but non-metric dissimilarity matrix, $M$, which serves as the ground truth for our evaluations. Then, we randomly removed 30% of the pairwise dissimilarities to introduce missing values and used the resulting matrix ($D$) as the input of the algorithms.
@@
    Multi-Scale Analysis: For a fixed sparsity level (30% missing), we generated three variants of $D$ with $m=25,50,$ and $100$ objects.
@@ 3.1.3 Evaluation Metrics
Performance was assessed by comparing the distances in the resulting Euclidean embedding, $\hat{D}$, to the complete, ground-truth non-metric matrix, $M$.
@@ 3.2 Data Characteristics
A critical first step was to quantify the non-metric nature of our synthetic data. By computing the Gram matrix and its eigenvalues, we can measure the deviation from Euclidean geometry. The non-metric Deviation Score, calculated as the ratio of the sum of the absolute values of negative eigenvalues to the sum of positive eigenvalues (Gower,, 1985), provides a standardized measure of this distortion. For our primary test case ($m=50$, 30% missing), the deviation score was 0.52, with the negative eigenvalues accounting for over a third of the total variance. This confirms the dataset is non-metric, presenting a significant challenge for any embedding algorithm.
@@ 3.3.1 Accuracy and Distance Reconstruction
In the main test case ($m=50$, 30% sparsity), Topolow demonstrated substantially higher accuracy in reconstructing the ground-truth dissimilarities compared to both Classical and Iterative MDS. As shown in Table 1, Topolow achieved a mean Normalized Stress of $0.190\pm 0.0003$, an improvement over Classical MDS (0.602) and Iterative MDS (0.552).
Table 1: Performance comparison on a non-metric dataset of 50 objects with 30% missing values. Values for stochastic methods are mean $\pm$ s.d. over 50 runs. Lower stress is better; higher $R^{2}$ is better.
@@
Figure 1: Shepard plots comparing embedded Euclidean distances to original non-metric dissimilarities for the best run of each method ($m=50$, 30% missing). The identity line (dashed) represents perfect reconstruction. Topolow’s embedding (right) shows a strong linear relationship and low distortion. Classical MDS (left) and Iterative MDS (center) show substantially more scatter and deviation from the ideal.
@@ 3.3.2 Robustness to Data Sparsity
At 90% sparsity—an extreme condition where only 10% of dissimilarities are known—Topolow achieved a mean Normalized Stress of $0.510$. This result is particularly striking as it is better than the performance of both MDS methods at the much lower 30% sparsity level. This demonstrates the power of Topolow’s approach: the interconnected system of springs enforces a global geometric consistency that allows for robust inference of the overall structure, even when the vast majority of direct measurements are absent. This is fundamentally more effective than attempting to embed a matrix where 90% of the values have been imputed.
@@ 4.1 The Input Dissimilarity Space
The algorithm’s mechanics treat each $D_{ij}$ as an independent target for the spring rest length between particles $i$ and $j$. If $D_{ij}\neq D_{ji}$, the algorithm will simply use the value corresponding to the specific pair $(i,j)$ being considered in a given step. Violations of the triangle inequality are resolved by the system of forces; if $D_{ik}>D_{ij}+D_{jk}$, the springs for $(i,j)$ and $(j,k)$ will pull those particles closer, while the spring for $(i,k)$ pulls them apart. The final configuration $X$ represents a minimum-energy compromise that best fits all constraints simultaneously, effectively finding a Euclidean projection of the non-metric input. ∎
@@ 5 Conclusion
We have presented the mathematics and evaluations of Topolow, an algorithm for Euclidean embedding of potentially sparse, noisy, and non-metric dissimilarity data. By adapting concepts from force-directed graph drawing for the purpose of quantitative metric reconstruction and employing a gradient-free, sequential pairwise optimization, Topolow overcomes the primary limitations of MDS methods. Its foundation in a Laplace error model provides robustness to outliers and censored data, and its integrated, likelihood-based approach to dimensionality selection removes the need for ad-hoc choices. Our empirical evaluation on severely non-metric synthetic data demonstrates that Topolow consistently produces more accurate and stable embeddings than standard MDS methods, particularly as data sparsity and problem size increase. This theoretical robustness translates to practical advantages, as shown in its original application (Arhami and Rohani,, 2025), where Topolow achieved comparable accuracy to MDS for well-structured data of antigenic dissimilarities of influenza viruses but demonstrated a 56% and 41% improvement for more complex dengue and HIV data, respectively. Furthermore, it exhibited orders of magnitude greater run-to-run stability and, unlike MDS, consistently produced a complete embedding of all viruses, a critical feature for sparse datasets. The algorithm’s ability to operate on general, non-metric dissimilarity data is one of its most critical features, providing a necessary bridge between raw, real-world measurements and the vast suite of analytical tools that require metric inputs. This makes Topolow a powerful and widely applicable tool for data analysis, visualization, and dimensionality reduction.